Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 109; σ = 14

P(x ≥ 120) =
Answer #1

z score for 120 = (120 - 109)/14 = 0.79

Hence,

P(x > 120)

= P(z > 0.79)

= 0.2160